If a 1300 kg car can accelerate from 30 km/h to 65 km/h in 3.6 s , how long will it take to accelerate from 65 km/h to 90 km/h ? Assume the power stays the same, and neglect frictional losses.
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2019-03-29T11:16:52-0400

Let's first convert km/h to m/s:


"v_f = 65 \dfrac{km}{h} \cdot \dfrac{1000m}{1 km} \cdot \dfrac{1 h}{3600 s} = 18.05 \dfrac{m}{s},"

"v_i = 30 \dfrac{km}{h} \cdot \dfrac{1000m}{1 km} \cdot \dfrac{1 h}{3600 s} = 18.05 \dfrac{m}{s}."

By the definition of the power, we get:


"P = \dfrac{W}{t},"

here, "W" is the work done by the car and "t" is time.

Let's find the work done by the car from the Work-Kinetic Energy theorem. It states that the work done by the car is equal to the change in kinetic energy of the car:


"W = \Delta KE = KE_f - KE_i,""W = \dfrac{1}{2}mv_f^2 - \dfrac{1}{2}mv_i^2 = \dfrac{1}{2}m(v_f^2 - v_i^2)."

Substituting "W" into the first formula we can find the power of the car when it accelerates from 30 km/h to 65 km/h in 3.6 s:


"P = \dfrac{\dfrac{1}{2} \cdot 1300kg \cdot((18.05 \dfrac{m}{s})^2 - (8.3 \dfrac{m}{s})^2)}{3.6 s} = 46387 W."

Now, let's find the work done by the car when it accelerates from 65 km/h to 90 km/h from the same Work-Kinetic Energy theorem, but, first, convert km/h to m/s:


"v_f = 90 \dfrac{km}{h} \cdot \dfrac{1000m}{1 km} \cdot \dfrac{1 h}{3600 s} = 25 \dfrac{m}{s},""v_i = 65 \dfrac{km}{h} \cdot \dfrac{1000m}{1 km} \cdot \dfrac{1 h}{3600 s} = 18.05 \dfrac{m}{s},""W = \dfrac{1}{2} \cdot 1300 kg \cdot ((25 \dfrac{m}{s})^2 - (18.05 \dfrac{m}{s})^2) = 194478 J."

Since, the power stays the same we can find the time that the car takes to accelerate from 65 km/h to 90 km/h from the same formula for the power:


"P = \dfrac{W}{t},""t = \dfrac{W}{P} = \dfrac{194478 J}{46387 W} = 4.2 s."

Answer:

"t = 4.2 s."